OpenStudy (anonymous):

how do you do f(x)=-2x+3 and g(x)=4x-3,which is greater,f(5) or g(-2)

OpenStudy (anonymous):

f(5)=-10+3=-7 g(-2)=-8-3=-11 f(5) is greater